# Deep Joint CSI Estimation-Feedback-Precoding for MU-MIMO OFDM Systems

### v1

@@ IV-E Scalability Tests
Figure 10: Testing the downlink spectral efficiency of CSI estimation, feedback, and precoding design networks under partially separate and joint architectures as affected by the amount of feedback.
Figure 11: Testing the scalability of the network architecture with a larger number of users.
@@ V Conclusion
In this paper, we propose an end-to-end architecture for CSI acquisition and precoding design in MU MIMO-OFDM systems, called JEFPNet. At the UE side, a joint CSI estimation and compression network is proposed to compress and extract only the semantic information relevant to precoding from received pilot signals, rather than the entire CSI. This approach eliminates the overhead of explicit CE and indirectly reduces the source compression rate to minimize information loss. At the BS, a proposed joint MU CSI reconstruction and precoding design network directly designs the precoding vectors from the received semantic information fed from each user. Additionally, a pilot learning matrix is employed at the BS to enhance the extraction efficiency of semantic information and maximize its utilization. The entire network is trained end-to-end with the objective of maximizing the downlink spectral efficiency. Multiple modules are jointly optimized to ensure that only semantic information related to the optimization objective is extracted and transmitted throughout the transmission process. Experimental results demonstrate that the proposed network achieves a higher downlink spectral efficiency under the same spectrum resource overhead. Furthermore, JEFPNet significantly reduces both storage overhead and computational complexity at the UE side, which is more suitable for UE side without powerful storage capacity and computational ability. Simultaneously, it decreases computational complexity at the BS, reduces precoding design delay, and equips the network with the ability to counteract channel aging compared to separate architecture. Finally, we evaluate the scalability of the proposed network under varying feedback overheads and numbers of users, demonstrating that the performance of the network under the proposed joint architecture consistently outperforms that of the separate architecture.

### v2

@@ IV-E Analysis of Storage and Computational Complexity Overhead
Figure 14: Testing the downlink spectral efficiency of joint CE-feedback-precoding networks under partially separate and joint architectures as affected by the amount of feedback.
IV-F Scalability Tests
In this subsection, we firstly experimentally verify the scalability of the proposed network under varying feedback overheads. First, we fix the $\mathrm{SNR_{ce}}$ of the CE phase to 10 dB, and the CE overhead ($L$) to 6. The maximum number of user $K_{\max}$ set to 6, and the evaluated number of user $K$ is 2. We compare the downlink spectral efficiencies of different network architectures with single-user CSI feedback overhead ($Z$) values of 12, 16, and 32. The experimental results, shown in Fig. 14, indicate that as feedback overhead increases, the downlink spectral efficiency for both joint and separate architectures gradually improves. As the feedback bandwidth increases, the performance gap between the two methods gradually narrows. Specifically, the downlink spectral efficiency of JEFPNet exceeds that of SEFPNet by 0.416 bps/Hz when $Z=12$, 0.387 bps/Hz when $Z=16$, and 0.333 bps/Hz when $Z=32$, under $\mathrm{SNR_{u}}=-10$ dB.
Figure 15: CSI rate-efficiency bounds.
In this paper, we adopt the DJSCC network architecture in the analog domain. To facilitate deployment in existing communication systems and maintain compatibility with current hardware, we assess the rate–efficiency bound of the proposed network in a digital communication system, analogous to the CSI rate–distortion bounds presented in [rate_distortion_1, rate_distortion_2]. Specifically, the downlink CSI features compressed at the UE side are mapped to the nearest power-normalized quadrature amplitude modulation (QAM) constellation points for transmission, while the BS utilizes the received constellation points for multiuser precoding design. In this part of the simulation experiments, we consider a CE phase with $\mathrm{SNR_{ce}}=10$ dB, a pilot length $L=6$, and modulation schemes including 4-QAM, 16-QAM, 64-QAM, 256-QAM, and 1024-QAM. These correspond to symbol quantization levels of 2, 4, 6, 8, and 10 bits, respectively. In this way, the feedback bandwidth $Z$ is maintained at a constant value of 32. The corresponding simulation results are presented in Fig. 15.
The simulation results in Fig. 15 demonstrate that the performance of JEFPNet in the digital domain gradually converges to that in the analog domain as the number of quantization bits increases, regardless of the number of users or uplink feedback $\mathrm{SNR_{u}}$. For instance, under the condition $K=2$, $\mathrm{SNR_{u}}=10$ dB, and 1024-QAM, the performance loss is only 0.048 bps/Hz. Similarly, with $K=6$ under the same conditions, the loss is limited to 0.093 bps/Hz. This improvement arises because increasing the number of quantization bits leads to a higher modulation order, resulting in a denser constellation and thereby reducing the loss and error introduced by mapping the analog signal to the constellation points.
IV-G Scenario Shift Evaluation
TABLE III: Scenario shift testing in UMi and UMa environments with downlink spectral efficiency.
bps/Hz $\mathrm{SNR_{u}}$ $\mathrm{-10}$ dB $\mathrm{0}$ dB $\mathrm{10}$ dB
$K$ Test Train UMa UMi UMa UMi UMa UMi
2 UMa 9.41 9.36 9.78 9.74 9.81 9.78
UMi 9.57 9.47 10.33 10.18 10.41 10.26
4 UMa 12.52 12.36 13.12 12.96 13.18 13.02
UMi 12.57 12.48 13.89 13.76 14.04 13.91
Investigating the scenario-shift problem faced by AI models is critical for next-generation communication systems that integrate AI with communication technologies, and it has been identified as a key research challenge in the 6G standardization process by 3GPP [3gpp38843]. Accordingly, in this set of experiments, we evaluate the impact of scenario shift on the performance of the proposed network. Specifically, we introduce additional Urban Micro (UMi) scenarios in comparison to the previously considered UMa scenarios. The UMi dataset is also generated using QuaDriGa following 3GPP TR 38.901 and TR 38.873. Except for scenario-specific parameters, such as the BS height, all other settings remain consistent with those of the UMa dataset to ensure fairness in comparison. We train JEFPNet under both UMa and UMi scenarios and subsequently test the trained models in both environments to assess the impact of scenario shifts on network performance, using downlink spectral efficiency as the evaluation metric. In this part of the experiment, we set $K_{\mathrm{max}}=4$, $Z=32$, $L=6$, and ${\mathrm{SNR_{ce}}}=10$ dB. We then evaluated the impact of scenario shift on JEFPNet performance in both UMi and UMa scenarios, considering $\mathrm{SNR_{u}}\in\{-10~\text{dB},0~\text{dB},10~\text{dB}\}$ and $K\in\{2,4\}$. The experimental results are presented in Table. III.
The experimental results indicate that, without scenario shifts, performance in the UMi scenario is better than in the UMa scenario. For example, with $K=2$ and $\mathrm{SNR_{u}}=-10$ dB, JEFPNet in the UMi scenario outperforms that in the UMa scenario by 0.16 bps/Hz. Similarly, with $K=6$ and $\mathrm{SNR_{u}}=10$ dB, the performance improvement reaches 0.86 bps/Hz. This advantage arises because the spatial characteristics of the UMi scenario are more favorable than those of the UMa scenario, enabling convergence to a better solution under identical training conditions. A similar phenomenon was observed in [UMi_UMa], where the CE error of the UMi scenario is smaller than that of the UMa scenario under the same conditions. In the presence of scenario shifts, the network trained in the UMa scenario performs worse than those trained in the UMi scenario when evaluated in the UMi scenario, although the performance gap narrows at lower $\mathrm{SNR_{u}}$. For instance, at $K=2$ and $\mathrm{SNR_{u}}=-10$ dB, the gap is only 0.16 bps/Hz. Conversely, networks trained in the UMi scenario often achieve superior performance than those trained in the UMa scenarios when tested in the UMa scenario, despite a slight decline, owing to their more favorable convergence positions.
@@ V Conclusion
In this paper, considering the channel coding and feedback channel, we propose an end-to-end architecture for CSI acquisition and precoding design in MU MIMO-OFDM systems with user-number adaptability, called JEFPNet. At the UE side, a joint CE and compression network is proposed to compress and extract only the semantic information relevant to precoding from received pilot signals, rather than the entire CSI. This approach eliminates the overhead of explicit CE and minimize information loss. At the BS, a proposed joint MU CSI reconstruction and Transformer-based precoding design network directly designs the precoding vectors from the received semantic information fed from each user. By leveraging a multi-head attention mechanism, the precoding design module can be adapted to varying numbers of users without substantially increasing the BS parameter count, thereby reducing model storage requirements, distribution overhead, and switching latency. Additionally, a pilot learning matrix is employed at the BS to enhance the extraction efficiency of semantic information and maximize its utilization. The entire network is trained end-to-end with the objective of maximizing the downlink spectral efficiency. Multiple modules are jointly optimized to ensure that only semantic information related to the optimization objective is extracted and transmitted throughout the transmission process.
Experimental results demonstrate that the proposed network achieves a higher downlink spectral efficiency under the same spectrum resource overhead. Furthermore, JEFPNet significantly reduces both storage overhead and computational complexity at the UE side, which is more suitable for UE side without powerful storage capacity and computational ability. Finally, we evaluate the scalability of the proposed network under varying feedback overheads and its adaptability to different numbers of users, demonstrating that the performance of the network under the proposed joint architecture consistently outperforms that of the separate architecture.